Jon can mow the lawn in 6 hours. Kim can mow the lawn in 5 hours. How long will it take them to mow the lawn if the work together?
Answer by nerdybill(7384)

Jon can mow the lawn in 6 hours. Kim can mow the lawn in 5 hours. How long will it take them to mow the lawn if the work together?
.
Let x = time (hours) it takes for both
then
x(1/6 + 1/5) = 1
multiplying both sides by 30:
x(5 + 6) = 30
11x = 30
x = 30/11
x = 2.73 hours
or
x = 2 hours and 44 minutes